The electric potential difference between the ground and a cloud in a particular 

thunderstorm is 1.2X10^9 V. In the unit of electron- volt, what is the magnitude of the 

electric potential energy of an electron that moves between ground and the clouds?



1
Expert's answer
2021-11-14T17:16:25-0500

"\Delta U_{GC}=e\cdot\Delta V\ (in \ J)"


"\Delta U_{GC}=e\cdot\Delta V/e=\Delta V\ (in \ eV)"


"\Delta U_{GC}=\Delta V=1.2\cdot10^9\ (eV)=1.2\ (GeV)" . Answer